OpenStudy (anonymous):

I need help with this problem. I have no idea how to go about solving it. A 10 inch diameter saw blade spins at 2000 revolutions per minute. What is the angular speed in radians per second? What is the linear speed in feet per second?

OpenStudy (amistre64):

how many radians are in 1 revolution?

OpenStudy (anonymous):

2pi.

OpenStudy (amistre64):

good, then there are 2pi*2000 revolutions per minute how many seconds in a minute?

OpenStudy (anonymous):

60 seconds in a minute.

OpenStudy (amistre64):

correct, so lets just do this \[\frac{2\pi~(2000)~rads}{1~min}\] \[\frac{2\pi~(2000)~rads}{60~sec}\]

OpenStudy (anonymous):

Do I multiply them together, or separately?

OpenStudy (loser66):



OpenStudy (amistre64):

i just replaced 1 min byt eh equaivalent 60 secs; then its a simple reduction

OpenStudy (amistre64):

my physics teacher would get mad if i didnt do it his way but he got over it when I kept getting the highest grade in the class lol

OpenStudy (anonymous):

Okay, so let me make sure I'm doing this correctly - I set my calculator to radian mode, then work out the equation you gave?

OpenStudy (amistre64):

radian mode is not needed ... we are not defining any sin cos trig function

OpenStudy (anonymous):

I got 200pi/3

OpenStudy (amistre64):

since 1 full revolution of a circle is equal to 2 pi radians then 2000 revolutions will be equal to 2pi * 2000 radians

OpenStudy (loser66):

@amistre64 aha!! now I know who I should tag on my physics problem.

OpenStudy (amistre64):

200pi/3 is good; that is the number of radians it is spinning in 1 second

OpenStudy (amistre64):

since 1 radian measures an arc of 1 radius; the linear calculation is 200pi/3 * (radius)